# From Acceleration to Saturation: Scaling Behavior of Bootstrapped Language Model Pretraining

### v1

@@ Abstract
Bootstrapped pretraining, i.e., the reuse of a pretrained base model for further pretraining, such as continual pretraining or model growth, is promising at reducing the cost of training language models from scratch. However, its effectiveness remains unclear, especially when applied to overtrained base models. In this work, we empirically study the scaling behavior of bootstrapped pretraining and find that its scaling efficiency diminishes in a predictable manner: The scaling exponent with respect to second-stage pretraining tokens decreases logarithmically with the number of tokens used to pretrain the base model. The joint dependence on first- and second-stage tokens is accurately modeled by a simple scaling law. Such saturation effect reveals a fundamental trade-off in multi-stage pretraining strategies: the more extensively a model is pretrained, the less additional benefit bootstrapping provides. Our findings provide practical insights for efficient language model training and raise important considerations for the reuse of overtrained models.
@@ 1 Introduction
Large language models (LLMs) have recently shown astounding performance in various natural language processing tasks, reaching human-level capabilities in some cases achiam2023gpt ; claude3 ; geminiteam2023gemini . However, training/pretraining these models from scratch is computationally expensive and time-consuming, requiring weeks to months even with powerful GPU clusters. To address this challenge within the pretraining stage itself, researchers have explored strategies for bootstrapping/reusing existing pretrained (base) models for various purposes. These include strategies to learn with new pretraining data, or to increase the model size, without starting from scratch. We refer such strategies that pretrain a model in multiple stages collectively as bootstrapped pretraining.
@@
More specifically, we are interested in bootstrapped pretraining strategies that (1) improve domain-specific performance through continual pretraining (CPT) of the base model (see, e.g., ibrahim2024simple ); or (2) scale model capacity via model growth techniques, which increase the model size reusing base model parameters to speed up training chen2015net2net . These strategies have been shown to be promising at accelerating performance improvement of LLMs while reducing the computational cost compared to pretraining from scratch.
On the other hand, neural networks could lose plasticity after training for a long time, thereby reducing their ability to learn from new data effectively berariu2021study ; ash2020warm ; sodhani2020toward . Grown models may also similarly suffer from bad initialization, resulting in ineffective utilization of their enlarged capacity. Hence, it is unclear if the second stage of pretraining can be effectively applied to models that are overtrained, or whether the scaling behavior remains consistent across different training stages and model sizes.
@@
In this paper, we seek to understand the scaling behavior of bootstrapped pretraining methods by running a multitude of controlled language modeling experiments. We find that, from the performance (cross-entropy loss) perspective, bootstrapping overtrained base models leads to saturation in scaling during the second-stage pretraining. Specifically, we show that this effect is quantifiable via scaling laws: The scaling exponent (with respect to the number of training tokens in the second stage) decreases as the base model is trained for longer periods of time. The decrease is proportional to the logarithm of the number of training tokens invested to the first stage. See Figure 1.
@@
Figure 1: Bootstrapped pretraining with overtrained base models leads to saturation in scaling behavior. Left: $D_{2}$ has power-law scaling. We show scaling behavior of second-stage training tokens ($D_{2}$) for different values of first-stage tokens ($D_{1}$). Middle: Interaction term explains decreasing exponents. The fitted exponents in the left plots are used to fit Equation 5 as a function of $D_{1}$, and are shown to agree well with the functional form. Right: Scaling factor has power-law scaling w.r.t. $D_{1}$. Top: Continual pretraining (CPT) on code data. Bottom: Model growth by stacking.
This saturation effect underscores a fundamental trade-off in multi-stage pretraining strategies: bootstrapping a more heavily pretrained base model does not necessarily lead to better performance. Our scaling laws enable a quantitative assessment of when bootstrapping is beneficial versus when training from scratch is preferable, offering practical guidance for efficient language model training.
Contributions. More technically, our main contributions are summarized as follows:
@@
  • We conduct extensive experiments on language models of various sizes and pretraining tokens/datasets, to study the scaling behaviors of bootstrapped pretraining methods, including continual pretraining and model growth.
@@
  • We formulate and empirically validate possible functional forms to quantify the the scaling behavior of these methods. We find that for a wide range of configuration of these methods, the following scaling relation holds:
    $L(D_{1},D_{2})=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E$
@@
    where $L$ is the validation loss after the second stage, $D_{1}$ and $D_{2}$ are the number of training tokens in the first and second stages, respectively, and $A,\alpha_{1},\alpha_{2},\alpha_{3},E$ are positive constants. We denote the term $\alpha_{3}\log D_{1}$ as the interaction term. The scaling exponent with respect to $D_{2}$ then quantifies the saturation effect of bootstrapped pretraining.
  • We also formulate a joint scaling law incorporating model size in addition to dataset sizes, see Equation 7. Using these results, we discuss the practical implications of the scaling laws for efficient bootstrapping of language models. Particularly, we show how the scaling laws provide guidance on when to train from scratch instead of bootstrapping, and how compute optimality is affected by bootstrapped pretraining.
@@ 1.1 Related Work
Continual pretraining. CPT has been widely used to adapt existing LLMs to specific domains sun2020understanding ; jang2022temporal ; jang2022continual , including code yadav2023continualcode ; zan2022cert ; guo2024deepseek and mathematics gong2022continualpretraining ; shao2024deepseekmath domains, to be studied in this work. Systematic studies at scale are relatively fewer, and our methodology mainly follows ibrahim2024simple .
@@
Model growth. While chen2015net2net was the first model growth work in the deep learning era, the idea of growing neural networks from smaller ones can be traced to the 90s NIPS1989_69adc1e1 ; Fahlman1990TheRC . More recent language model-related model growth methods include gong2019efficient ; chen2021bert2bert ; wang2023learning ; wang2023lemon ; shen2022staged ; evci2022gradmax ; yao2024masked , which were subsequently systematically analyzed in du2024stacking , motivating our choice of model growth methods.
Scaling laws. We focus on modeling only the final loss, which has a power-law like behavior, a generic phenomenon not only occuring in neural networks hestness2017deep ; kaplan2020scaling , but is also observed in other natural as well as man-made phenomena clauset2009power . We shall additionally note that recent scaling law studies attempted to fit the whole loss curve albeit with more sophisticated functional forms tissue2024scaling ; wang2025learning ; qiu2025scaling .
@@
Similar to CPT, the reduced capability of transfer learning of code data from models pretrained on natural language data was observed in hernandez2021scaling , where the authors denoted as ossification. A similar phenomenon was also observed in the pretraining-instruction-tuning pipeline springer2025overtrained . These studies however did not quantify the saturated scaling behavior. Previous fine-tuning scaling studies mikami2022scaling ; zhang2024scaling ; bethune2025scaling did not observe the subtler saturation effects as well, which is perhaps due to the smaller scale (in terms of training tokens) of their experiments.
2 Experimental Setup
@@
Model configuration. We consider decoder-only transformers vaswani2017attention pretrained with an autoregressive language modeling objective. Our architecture is LLaMA-like touvron2023llama , incorporating refinements such as SwiGLU activation functions shazeer2020glu and rotary position embeddings su2024roformer . We use the LLaMA tokenizer with a vocabulary size of 32,000. We consider a suite of model sizes up to 1.1B in our experiments. Table 3 of Appendix B contains the model configuration used in this paper.
@@
Training configuration. All experiments are conducted using the Megatron-LM library shoeybi2019megatron . We train models in mixed precision (bfloat16) using the AdamW optimizer loshchilov2017fixing , with maximum learning rates tuned separately for each model size. The learning rate follows the warmup-stable-decay (WSD) schedule bi2024deepseek ; hu2024minicpm , with the final learning rate decayed to one tenth of the peak value. This schedule enables long training runs and facilitates checkpoint reuse for emulating shorter effective training budgets, reducing computational cost. In Appendix C, we also provide an ablation study with the cosine learning rate schedule to show that they achieve similar performance. We train each model for a number of tokens for up to roughly 200 times the model size in tokens in total (up to a few hundred B tokens overall). Tables 4 and 5 in Appendix summarize the main training configurations. See also Appendix B for further experimental details and Appendix C for training details and evaluation on downstream tasks.
Two stages of pretraining.
@@
In the context of bootstrapped pretraining, we are primarily interested in how the loss behaves with respect to the number of training tokens used in the two stages of pretraining:
@@
For the first-stage pretraining, we use the CommonCrawl portion of the Slimpajama-DC dataset shen2023slimpajama , containing 368B tokens in total. The models trained in this stage serve as base checkpoints for the second-stage bootstrapped pretraining experiments.
Continual Pretraining.
@@
CPT refers to further training a pretrained model on a large dataset, typically billions of tokens, with the goal of improving performance on a new domain. We distinguish CPT from instruction tuning, which typically trains with smaller datasets (millions of tokens or fewer), and is thus harder to analyze the scaling behavior. For our experiments, we perform CPT on the base model on two domain-specific datasets: Code corpus: Stack/StarCoder li2023starcoder ; and mathematics corpus: OpenWebMath paster2023openwebmath . Unless otherwise noted, we adopt the same optimizer and learning rate configurations as the first-stage pretraining.
Model growth.
@@
Model growth refers to increasing model capacity by adding new layers and/or expanding hidden dimensions, thereby increasing the number of trainable parameters. This strategy aims to leverage the representations learned during first-stage training, allowing the larger model to accelerate learning in the second stage. We investigate two model growth techniques shown to be the most effective (in terms of adding new parameters in the width and depth directions) in prior work du2024stacking (see also Appendix C.5 for more details):
@@
  • Width expansion: Add new neurons to each layer while using function-preserving initialization (FPI), ensuring that the expanded model initially reproduces the behavior of the smaller model chen2021bert2bert ; du2024stacking .
@@
  • Depth-wise stacking: Add new layers to the top of the transformer by copying existing layer weights, thereby extending model depth chen2021bert2bert ; du2024stacking .
@@
After performing this procedure, the larger models are trained on the same dataset as the base model. For both bootstrapped pretraining scenarios in consideration, the validation loss is evaluated on a held-out set from the second-stage dataset. We illustrate the whole framework in Figure 2.
Figure 2: Illustration of bootstrapped pretraining in consideration. Bootstrapped pretraining consists of two stages: (1) first-stage pretraining of a base model for $D_{1}$ tokens on internet/generic data; (2) bootstrapping/second-stage pretraining via continual pretraining or model growth for $D_{2}$ tokens. Section 3 and 4 study and develop scaling laws as a function of these two variables (and additionally model size, NN in Section 5) to predict the final loss after the second stage.
3 Formulating the Data Scaling Laws
In this Section, we derive possible functional forms for the scaling laws of bootstrapped pretraining methods, focusing on the number of training tokens as the main variable of interest. These are then fitted and compared empirically in later sections.
Power-law ansatz.
We begin by assuming that the validation loss $L$ follows the widely observed power-law scaling with respect to a single variable of interest, such as the number of training tokens or the model size hestness2017deep ; hestness2019beyond ; henighan2020scaling :
$L=AX^{-\alpha}+E,$
where $X$ is the variable of interest, $\alpha$ is the scaling exponent, $A$ is the scaling factor, and $E$ is the irreducible loss due to the inherent entropy of the data distribution.  [^1] Strictly speaking, we assume the form $L=\tfrac{A}{(X+1)^{\alpha}}+E$, which ensures finiteness at $X\to 0$. However, since in practice $X\gg 1$ (typically $10^{6}$ or more), we approximate it as $L\approx AX^{-\alpha}+E$ for notational simplicity. Moreover, we use the same symbols $A$, $\alpha$, and $E$ to denote the scaling factor, exponent, and additive constant, respectively, for notational conveniences, although they may differ across different variables of interest.

### v2

# Reusing Overtrained Language Models Saturates Scaling

@@ Abstract
Reusing pretrained base models for further pretraining, such as continual pretraining or model growth, is promising at reducing the cost of training language models from scratch. However, the effectiveness remains unclear, especially when applied to overtrained base models. In this work, we empirically study the scaling properties of model reuse and find that the scaling efficiency diminishes in a predictable manner: The scaling exponent with respect to second-stage training tokens decreases logarithmically with the number of tokens used to pretrain the base model. The joint dependence on first- and second-stage tokens is accurately modeled by a simple scaling law. Such saturation effect reveals a fundamental trade-off in multi-stage pretraining strategies: the more extensively a base model is pretrained, the less benefit additional pretraining provides. Our findings provide practical insights for efficient language model training and raise important considerations for the reuse of overtrained models.
language modeling, scaling law
@@ 1 Introduction
Large language models (LLMs) have recently shown astounding performance in various natural language processing tasks, reaching human-level capabilities in some cases (Achiam et al., 2023; Anthropic, 2024; Gemini Team Google, 2023). However, training/pretraining these models from scratch is computationally expensive and time-consuming, requiring weeks to months even with powerful GPU clusters. To address this challenge within the pretraining stage itself, researchers have explored strategies for reusing existing pretrained (base) models for various purposes. These include strategies to learn with new pretraining data, or to increase the model size, without starting from scratch. Such strategies, which pretrain a model through multiple stages, are collectively referred to as model reuse throughout this paper.
@@
More specifically, we are interested in strategies that (1) improve domain-specific performance through continual pretraining (CPT) of the base model (see, e.g., Ibrahim et al. (2024)); or (2) scale model capacity via model growth techniques, which increase the model size reusing base model parameters to speed up training (Chen et al., 2015). These strategies have been shown to be promising at accelerating performance improvement of language models while reducing computational costs of pretraining from scratch.
On the other hand, when a base model is overtrained, intuitively, its highly optimized parameters make effective exploration after capacity enlargement for model growth (or exploration with new data distribution for CPT) difficult, causing the second stage to learn slower. This is closely related to the loss of plasticity for neural networks over changing tasks (Ash and Adams, 2020; Sodhani et al., 2020; Lyle et al., 2023, 2024). Hence, it is unclear if the second stage of pretraining can be effectively applied to language models that are overtrained.
@@
In this paper, we seek to understand the scaling properties of such methods by running a multitude of controlled language modeling experiments. We find that, from the performance (cross-entropy loss) perspective, reusing overtrained base models leads to saturation in scaling during the second-stage pretraining. Specifically, we show that this effect is quantifiable via scaling laws: The scaling exponent (with respect to the number of training tokens in the second stage) decreases as the base model is trained for longer periods of time. The decrease is proportional to the logarithm of the number of training tokens invested to the first stage. See Figure 1. Our work bears practical implications as even relatively tiny models are overtrained to trillions of tokens nowadays (Zhang et al., 2024b; Dubey et al., 2024), and model reuse strategies are widely adopted.
@@
Figure 1: Model reuse with overtrained base models leads to saturation in scaling behavior. Left: $D_{2}$ has power-law scaling. We show scaling behavior of second-stage training tokens ($D_{2}$) for different values of first-stage tokens ($D_{1}$), trained on a 0.1B base model. Middle: Interaction term explains decreasing exponents. The fitted exponents in the left plots are used to fit Equation 3 as a function of $D_{1}$, and are shown to agree well with the functional form. Right: Scaling factor has power-law scaling w.r.t. $D_{1}$. Top: Continual pretraining (CPT) on code data. Bottom: Model growth from 0.1B to 0.2B by stacking.
Summary of contributions. More technically, our key contributions are as follows:
@@
  • We conduct extensive experiments on language models of various sizes and pretraining tokens/datasets (with over 450 runs), to study the scaling properties of model reuse methods, including continual pretraining and model growth.
@@
  • We find that for a wide range of configuration of these methods, the following empirical scaling relation holds:
    $L(D_{1},D_{2})=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E$ (1)
@@
    where $L$ is the validation loss after the second stage, $D_{1}$ and $D_{2}$ are the number of training tokens in the first and second stages, respectively, and $A,\alpha_{1},\alpha_{2},\alpha_{3},E$ are positive constants. We denote the term $\alpha_{3}\log D_{1}$ as the interaction term. The scaling exponent with respect to $D_{2}$ then quantifies the saturation effect of model reuse. We also formulate other plausible functional forms to make comparisons, showing that the above formula still best fits the empirical observation.
  • We further analyze the scaling properties with respect to model size, and show that the scaling laws can be extended to jointly incorporate model size together with dataset sizes, see Equation 9. Limitations of our scaling law formulation when training with an extremely large number of tokens are also discussed.
  • We demonstrate how the scaling laws can provide guidance on when to train from scratch instead of model reuse to mitigate the negative effects of overtraining. Moreover, we analyze the saturation phenomenon both mechanistically and via a toy model.
@@ 1.1 Related Work
Continual pretraining. CPT has been widely used to adapt existing LLMs to specific domains (Sun et al., 2020; Jang et al., 2022a, b), including code (Yadav et al., 2023; Zan et al., 2022; Guo et al., 2024) and mathematics (Gong et al., 2022; Shao et al., 2024) domains, to be studied in this work. Systematic studies at scale are relatively fewer, and our methodology mainly follows Ibrahim et al. (2024).
@@
Model growth. While Chen et al. (2015) was the first model growth work in the deep learning era, the idea of growing neural networks from smaller ones can be traced to the 90s (Fahlman and Lebiere, 1989; Fahlman, 1990). More recent language model-related model growth methods include Gong et al. (2019); Chen et al. (2021); Wang et al. (2023a, b); Shen et al. (2022); Evci et al. (2022); Yao et al. (2024), which were subsequently systematically analyzed in Du et al. (2024), motivating our choice of model growth methods.
Power-law ansatz. We assume that the final validation loss $L$ follows the widely observed power-law scaling with respect to a single variable of interest, such as the number of training tokens or the model size (Hestness et al., 2017, 2019; Henighan et al., 2020):
$L=AX^{-\alpha}+E,$ (2)
where $X$ is the variable of interest, $\alpha$ is the scaling exponent, $A$ is the scaling factor, and $E$ is the irreducible loss due to the inherent entropy of the data distribution.  [^1] Strictly speaking, we assume the form $L=\tfrac{A}{(X+1)^{\alpha}}+E$, which ensures finiteness at $X\to 0$. However, since in practice $X\gg 1$ (typically $10^{6}$ or more), we approximate it as $L\approx AX^{-\alpha}+E$ for notational simplicity. Moreover, we use the same symbols $A$, $\alpha$, and $E$ to denote the scaling factor, exponent, and additive constant, respectively, for notational conveniences when there is no ambiguity, although they may differ across different variables of interest.
@@
Similar to CPT, the reduced capability of transfer learning of code data from models pretrained on natural language data was observed in Hernandez et al. (2021), where the authors denoted as ossification. A similar phenomenon was also observed in the pretraining-fine-tuning pipeline (Springer et al., 2025). These studies however did not quantify the saturated scaling behavior. Previous fine-tuning scaling studies (Mikami et al., 2022; Zhang et al., 2024a; Bethune et al., 2025) did not observe the subtler saturation effects as well, which is perhaps due to the smaller scale (in terms of training tokens) of their experiments.
2 Experiments
2.1 Setup
@@
Model configuration. We consider decoder-only transformers (Vaswani et al., 2017) pretrained with an autoregressive language modeling objective. Our architecture follows LLaMA (Touvron et al., 2023), incorporating refinements such as SwiGLU activation functions (Shazeer, 2020) and rotary position embeddings (Su et al., 2024). We use the LLaMA tokenizer with a vocabulary size of 32,000. Table 2 of Appendix B contains the model configuration used in this paper.
@@
Training configuration. All experiments are conducted using the Megatron-LM library (Shoeybi et al., 2019). We train models in mixed precision (bfloat16) using the AdamW optimizer (Loshchilov et al., 2017), with maximum learning rates tuned separately for each model size. The learning rate follows the warmup-stable-decay (WSD) schedule (Bi et al., 2024; Hu et al., 2024), with the final learning rate decayed to one tenth of the peak value. This schedule enables long training runs and facilitates checkpoint reuse for emulating shorter effective training budgets, reducing computational cost. In Appendix C, we also provide an ablation study with the cosine learning rate schedule to show that they achieve similar performance.
We train the models for a number of tokens to at least roughly 200 times the model size in tokens in total (token-to-parameter ratio). Tables 3 and 4 in Appendix summarize the main training configurations. See also Appendix B for further experimental details and Appendix C for training details and evaluation on downstream tasks.
Figure 2: Illustration of model reuse in consideration. model reuse consists of two stages: (1) first-stage pretraining of a base model for $D_{1}$ tokens on internet/generic data; (2) second-stage pretraining via continual pretraining or model growth for $D_{2}$ tokens. Section 3 and 3.2 study and develop scaling laws as a function of these two variables (and additionally model size, NN in Section 4) to predict the final loss after the second stage.
@@
Two stages of pretraining. In the context of model reuse, we are primarily interested in how the loss behaves with respect to the number of training tokens used in the two stages of pretraining:
@@
Base model. For the first-stage pretraining, we use the CommonCrawl portion of the Slimpajama-DC dataset (Shen et al., 2023), containing 368B tokens in total. The models trained in this stage serve as base checkpoints for the second-stage model reuse experiments.
@@
Continual Pretraining. CPT refers to further training a pretrained model on a large dataset, typically billions of tokens, with the goal of improving performance on a new domain. We distinguish CPT from instruction tuning, which typically trains with smaller datasets (millions of tokens or fewer), and is thus harder to analyze the scaling behavior. For our experiments, we perform CPT on the base model with two domain-specific datasets: Code corpus: Stack/StarCoder (Li et al., ); and mathematics corpus: OpenWebMath (Paster et al., 2023). Unless otherwise noted, we adopt the same optimizer and learning rate configurations as the first-stage pretraining.
@@
Model growth. Model growth refers to increasing model capacity by adding new layers and/or expanding hidden dimensions, thereby increasing the number of trainable parameters. This strategy aims to leverage the representations learned during first-stage training, allowing the larger model to accelerate learning in the second stage. We investigate two model growth techniques shown to be the most effective (in terms of adding new parameters in the width and depth directions) in prior work (Du et al., 2024) (see also Appendix C.5 for more details):
@@
  • Width expansion (exp): Add new neurons to each layer while using function-preserving initialization (FPI), ensuring that the expanded model initially reproduces the behavior of the smaller model.
@@
  • Depth-wise stacking (stk): Add new layers to the top of the existing one by direct copying, thereby extending model depth.
@@
After performing this procedure, the larger models are trained on the same dataset as the base model. The growth factor is defined as the ratio of the number of non-embedding parameters in the grown model to that in the base model. For both model reuse scenarios in consideration, the validation loss is evaluated on a held-out set from the second-stage dataset. We illustrate the whole framework in Figure 2.
2.2 Experimental Observations
We here provide experimental observations that motivate our formulation of the scaling law. In Figure 1, we show results of training a base model of size 0.1B, and performing CPT on code data, and stacking with a growth factor of 2 on a $5\times 5$ grid of $D_{1},D_{2}$.  [^2] We double the number of non-embedding layers for stacking. In the left panels, we plot the second-stage validation loss after the second stage as a function of $D_{2}$ for different values of $D_{1}$ (left). The following observation is made:
Observation 1: The loss is observed to follow a power law with respect to the number of tokens $D_{2}$, i.e., Equation 2.
This is consistent with previous scaling laws reflecting the expectation that, when starting from a fixed initialization, additional tokens improve performance predictably.
Furthermore, we observe that the fitted scaling exponent of Equation 2 decreases as $D_{1}$ increases, with the model growth method more pronouncedly so. To quantify this trend, we express (the minus of) the scaling exponent as a function of $D_{1}$, denoted by $\alpha(D_{1})$. A scatter plot of $\alpha(D_{1})$ reveals a clear logarithmic dependence:
$\alpha(D_{1})=\gamma\log D_{1}+E^{\prime}.$ (3)
This relationship fits the empirical data well, as shown in the same Figure. Substituting Equation 3 into Equation 2, we arrive at a term of the form $D_{2}^{-E^{\prime}+\gamma\log D_{1}}$. Additionally in the same Figure, we show that the multiplicative dependence, $A\propto D_{1}^{-\alpha_{1}}$, also holds well. These lead to the following observation:
Observation 2: The dependencies of the second-stage loss jointly on $D_{1}$ and $D_{2}$ can be captured by a multiplicative scaling law with an interaction term, i.e., Equation 1.
Interpreting the scaling law. Several insights can be made from the scaling law. Fixing $D_{1}$, the effective scaling factor for $D_{2}$ becomes $AD_{1}^{-\alpha_{1}}$. As $D_{1}$ increases, this factor decreases, resulting in a lower initial loss at the start of second-stage pretraining, agreeing with the conventional wisdom that better-pretrained base models (larger $D_{1}$) provide stronger initialization for second-stage pretraining.
3 Scaling Laws for Model Reuse
3.1 Functional Forms for the Scaling Laws
The observations from the previous Section regarding the functional form of the scaling law are purely empirical, which naturally raises the question of whether other functional forms might also fit the observed data. To explore these possibilities, we first derive functional forms that satisfy certain reasonable assumptions, and later fit and compare them empirically with our experimental results.